OpenStudy (anonymous):

Three consecutive integers that have a sum of 48

OpenStudy (anonymous):

Ooh that is a tricky problem. I will just have to think about it logically. What is 48/3? It is 16

OpenStudy (anonymous):

Knowing that, I know that I will need a sum that will equal 16(3) Well isn't 15+16+17 = 48 The 15 and 17 displace each other to make two 16 Does that make sense?

OpenStudy (anonymous):

hi :)

OpenStudy (anonymous):

if the middle number is \(n\), then the first number is \(n-1\) and the last number is \(n+1\). if you add \((n-1)+n+(n+1)\) you get \(3n\), so to find the middle number we merely solve:$$3n=48\\n=48/3=16$$ i.e. our middle number is \(16\). what are the first and last numbers?

OpenStudy (anonymous):

15 an 17 is the answer?

OpenStudy (anonymous):

the three numbers are 15, 16, and 17 -- yes!